# Tallahassee, FL 32306-4350, USA

### v1

@@ 1 Introduction
From a theoretical point of view, the problem of investigating the effects of unknown heavy NP lends itself naturally to an effective field theory (EFT) approach, where the effects of NP at the EW scale are encoded in non-renormalizable local interactions of the SM fields. The SM is then to be seen as the 4-dimensional part of a more general EFT whose Lagrangian also includes operators of higher dimension obtained from the projection of a given UV theory onto the space of fields of the SM. The SM Effective Field Theory (SMEFT) realizes this idea by extending the Lagrangian of the SM by a complete set of Lorentz and (SM-)gauge invariant effective interactions of increasing mass dimensionality, built in terms of SM fields. These interactions do, in general, break the SM accidental symmetries but could also satisfy additional discrete and global symmetry properties, depending on the kind of NP they encode (see Falkowski:2023hsg ; Isidori:2023pyp for a general introduction).
@@
Different UV extensions of the SM project differently on the effective interactions of the SMEFT Lagrangian below the UV scale $\Lambda$ deBlas:2017xtg ; Guedes:2023azv ; Guedes:2024vuf , such that constraining the effect of SMEFT interactions with available data could discriminate between classes of UV models and guide future BSM explorations. Most notably, different flavour symmetries assumed at the UV scale $\Lambda$ modify the number and structure of SMEFT interactions present below $\Lambda$ and directly affect a broad spectrum of observables. Several studies Aoude:2020dwv ; Faroughy:2020ina ; Bissmann:2020mfi ; Bruggisser:2021duo ; Bruggisser:2022rhb ; Bellafronte:2023amz ; Grunwald:2023nli ; Greljo:2023adz ; Allwicher:2023shc ; Bartocci:2023nvp ; Mantani:2025bqu have investigated the impact of different UV flavour assumptions and emphasised the need to move beyond the initial flavour-blind approach. A popular first step beyond assuming complete invariance under family rotation for the five kinds of fermions present in the SM (the so-called $U(3)^{5}$ flavour symmetry of the two left-handed doublets and three right-handed singlets of the SM gauge group) consists in singling out the third generation and still assuming invariance under rotation within the first two families (the so-called $U(2)^{5}$ flavour symmetry assumption, which describes to a good approximation the quark flavour structure of the SM). This choice is motivated by the scale hierarchy present among fermions and naturally appears in several BSM models. In our study we will consider both cases and discuss pros and cons of both assumptions in current fits.
@@
Overall, the great advantage of employing an effective field theory framework is that we can implement in a very general and systematic way a bottom-up approach to the exploration of some unknown UV theory using our knowledge of the physics we can currently test. Had patterns to emerge (e.g. enhanced or suppressed SMEFT coefficients, correlation and flat directions in SMEFT space, etc.) they could help in selecting classes of models whose investigation could then become more specific. The effective field theory framework is also useful in this regard. Using a top-down approach, one can easily re-interpret the SMEFT results in terms of specific models by using the matching between a given UV scenario and the effective theory. In the absence of unambiguous signs of BSM physics, the SMEFT general approach, although less informative, can indeed provide a very powerful exploratory tool. With this respect, we now have at our disposal for the first time a very broad spectrum of sufficiently accurate experimental measurements that can be simultaneously used to constrain the effective interactions of the SMEFT. This has indeed become a major effort in particle physics phenomenology and various comprehensive studies based on fits of the SMEFT that include a broad spectrum of observables have appeared in the recent literature Ellis:2020unq ; Ethier:2021bye ; Garosi:2023yxg ; Allwicher:2023shc ; Bartocci:2023nvp ; Celada:2024mcf ; ATLAS:2024lyh ; Bartocci:2024fmm ; CMS:2025ugn ; terHoeve:2025gey .
@@
We will be working in the SMEFT between the UV scale $\Lambda$ and the EW scale ($\mu_{W}$), while at $\mu_{W}$ we will transition to the Low Energy Effective Theory (LEFT) Jenkins:2017jig where all the heavy SM degrees of freedom ($W,Z,H,t$) are integrated out. To fully benefit of the EFT formalism and in order to account for impactful quantum effects, we will consider the scale dependence of the SMEFT and LEFT Wilson coefficients using leading-order RGE, with tree-level matching conditions at both the UV scale $\Lambda$ and the EW scale $\mu_{W}$. All fits are performed in the open-source HEPfit framework DeBlas:2019ehy ; hepfitsite using for each observable state-of-the-art theoretical predictions and the most recent experimental measurements.
The outcome of the fits presented in this paper will determine the ranges of SM parameters and SMEFT Wilson coefficients at the UV scale $\Lambda$ that are compatible with the selected broad set of experimental measurements. We will present both fits in which only one coefficient at a time is active and fits in which all SMEFT coefficients are simultaneously active at the scale $\Lambda$. Single-coefficient fits can be used to extract information on the set of observables that most constrain a given coefficient and can also give preliminary indications on the main RGE effects, while the global fit fully represents the combined effect of using a broad spectrum of observables and considering the operator mixing induced by the RGE. At the same time, a global fit of SM parameters and SMEFT coefficients can be very sensitive to the parametric and theoretical uncertaintes and thus less effective, given current accuracies, to constrain the theory at hand. This dilution of the sensitivity is less evident in the $U(3)^{5}$ scenario, but has a strong impact in the $U(2)^{5}$ case, calling for the inclusion of more observables and/or for higher theoretical and experimental precision on the currently included ones.
@@
The paper consists of this introductory section followed by three main core sections and two appendices. A final section summarises the main conclusions of our study and presents an outlook on future developments. Concerning the core sections, in Section 2 we review elements of the SMEFT relevant to our study with particular emphasis on the implementation of the two aforementioned flavour scenarios, namely $U(3)^{5}$ and $U(2)^{2}$, and a detailed discussion of how the RGE of the SMEFT coefficients and SM parameters is included in our study. Technical details on how these two scenarios are implemented in the SMEFT basis of operators considered in our study are provided in Appendix A. Section 3 is dedicated to a discussion of the fitting framework and procedure, including how the various sets of observables have been implemented in HEPfit. A complete set of results is illustrated and discussed in Section 4 while the full outcome of both individual and global fits, for all three UV scales Λ=1, $\Lambda=1,\,3,$ and 10 TeV, with and without RGE, are collected in a series of tables in Appendix B.
@@ 2.1 The Standard Model Effective Field Theory: formalism and assumptions
where ${\cal O}_{i}^{(d)}$ are Lorentz and (SM)-gauge invariant operators of canonical mass dimension $d$ built only of SM fields, $i$ collectively labels each operator according to its field composition and quantum numbers, and $\Lambda$ is the cutoff of the effective theory indicatively identified with the UV scale of NP. If lepton number is conserved by the NP, the leading-order effects appear at dimension 6. In this work, we truncate the $\mathcal{L}_{\mathrm{SMEFT}}$ to include only dimension-6 operators and adopt the Warsaw basis introduced in ref. Grzadkowski:2010es . We also neglect baryon number violating and CP-odd operators. This results in the set of operators listed in Table 1.
@@
Table 1: Dimension-6 operators in the Warsaw basis, adapted from ref. Grzadkowski:2010es . We collectively denote by $X$ the field strength tensors of the SM gauge group ($X=G^{\mu\nu},W^{\mu\nu},B^{\mu\nu}$), by $\psi$ a generic fermion field, and by $\phi$ the scalar field of the SM. Among the fermion fields, $q$ and $l$ denote $SU(2)_{L}$ left-handed doublets while $u,d,e$ $SU(2)_{L}$ right-handed singlets, with family indices specified by the superscript in square brackets.
Furthermore, we do not consider the operator $\mathcal{O}_{\phi}$ since it does not enter at leading order in any of the observables considered in this study. In fact, the main effect of this operator is to modify the Higgs-boson self-interactions, and can only be weakly constrained with current Higgs pair-production data CMS:2022dwd ; ATLAS:2024ish . We also notice that while this operator can be generated from other bosonic interactions via renormalization (to be discussed in Section 2.2), it does not enter in the 1-loop anomalous dimension matrix entries for any other SMEFT interaction, i.e. no other operator is generated from a non-zero value of $C_{\phi}$ at the UV scale.
@@
To define the $U(2)^{5}$ limit, it is necessary to specify a direction in flavour space that identifies the third generation. In this work, we choose the direction of the third family to be the one defined by the Yukawa couplings at the NP scale $\Lambda$, and consider the two possibilities where the third family is defined: 1) in the basis of diagonal up-type quark Yukawa couplings (denoted as UP basis in the following), or 2) in the basis of diagonal down-type quark Yukawa couplings (denoted as DOWN basis in the following). It is important to emphasise that these two classes of scenarios do not correspond to a mere change of basis, but rather reflect different assumptions about the flavour structure of the underlying UV physics. Notice that such a flavour misalignment arises exclusively in the left-handed quark sector — see the discussion in Silvestrini:2018dos — while right-handed quark Yukawa couplings, as well as the full charged-lepton Yukawa couplings, can always be chosen flavour diagonal.
Having specified the basis, the $U(2)^{5}$ limit is described by a total of 124 (123 if not considering $C_{\phi}$) independent Wilson coefficients Faroughy:2020ina . As in the $U(3)^{5}$ case, the relations between the $U(2)^{5}$ Wilson coefficients and those of the general SMEFT are presented in Appendix A.2, which also sets the notation for the presentation of the $U(2)^{5}$-symmetric results.
@@ 2.2 Renormalization group evolution in the SMEFT
Assuming that the SMEFT coefficients are generated by BSM physics at the UV scale $\Lambda$, the Wilson coefficients at the EW scale $\mu_{W}$ are obtained by running them from the UV scale $\Lambda$ to the EW scale $\mu_{W}$ using the leading order (LO) RG equations derived in refs. Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga . Lacking the next-to-leading SMEFT anomalous dimension, we can control only log-enhanced higher-order corrections to SMEFT contributions. Hence, we use tree-level matching conditions and matrix elements, and compute all SMEFT Wilson coefficients at a common scale $\mu_{W}$ that for convenience we choose to be $\mu_{W}=M_{W}$. For all high-$p_{T}$ observables (Higgs, top, and Drell-Yan), we use the same scale $\mu_{W}$, since the difference between the relevant physical scale of those processes and $M_{W}$ is such that no large logarithms arise, except for a few regions of phase space which currently still have large experimental uncertainties.
@@
One could be tempted to apply the same procedure to quark masses and CKM angles Descotes-Genon:2018foz . However, one of the symmetry assumptions we are using, namely $U(2)^{5}$, requires specifying the structure of the SM Yukawa couplings at the scale $\Lambda$, and the relations between quark masses, CKM angles, and Yukawa couplings are modified by SMEFT contributions. Thus, one cannot just trade the experimental values of quark masses and CKM angles for the corresponding SM parameters at the EW scale, and keeping the SMEFT contributions in the Yukawa matrices RGE is essential. [^1] To our knowledge, these corrections have been neglected in all the SMEFT fits that have considered the RGE evolution of the SMEFT coefficients so far. To impose the required flavour structure, we choose as parameters in the quark flavour sector the Yukawa couplings at the scale $\Lambda$, which are then evolved to the EW scale including the effect of the SMEFT in the running. Once evolved to the EW scale, the Yukawa couplings are used to compute the quark masses and CKM angles, including SMEFT contributions, which are then compared to the experimental values. Concerning the leptons, since we are assuming diagonal Yukawa couplings for the charged leptons and individual lepton numbers are preserved, the flavour orientation at the scale $\Lambda$ is irrelevant and, as stated above, we can trade the lepton Yukawa couplings at the EW scale for the corresponding masses.
The RG evolution from the scale $\Lambda$ to $\mu_{W}$ is performed within the HEPfit framework using the RGEsolver library DiNoi:2022ejg . The SMEFT Lagrangian at the scale $\mu_{W}$ is then expanded around the minimum of the scalar potential, and the quark mass matrices are diagonalised through the biunitary transformations
@@
The SMEFT operators involving quark fields are then transformed to the mass eigenstate basis using the matrices $V_{u,d}^{L,R}$ after removing unphysical phases. The unitary CKM matrix is defined as $V_{\mathrm{CKM}}\equiv V_{u}^{L\dagger}V_{d}^{L}$. [^2] Note that the determination of the CKM matrix elements from the experimental data is affected by additional corrections from SMEFT operators contributing to the relevant CKM observables. All these potential new physics effects are included in our calculations. The SMEFT coefficients in the mass eigenstate basis are then used to compute the observables at the scale $\mu_{W}$. [^3] Under our flavour assumptions, the flavour rotation is expected to have a negligible effect for the calculation of flavour-diagonal observables such as EW, Higgs boson, and top quark observables. Most observables in the flavour sector involve lower energies, and are therefore computed in the LEFT, where the massive gauge bosons, the Higgs field and the top quark are integrated out. The matching between the two EFT’s, the SMEFT and the LEFT, is implemented according to ref. Jenkins:2017jig . For each flavour observable, as detailed in Section 3.6, the LEFT coefficients are evolved to the relevant energy scale using the LEFT RG equations.

### v2

@@ 1 Introduction
From a theoretical point of view, the problem of investigating the effects of unknown heavy NP lends itself naturally to an effective field theory (EFT) approach, where the effects of NP at the EW scale are encoded in non-renormalizable local interactions of the SM fields. The SM is then to be seen as the 4-dimensional part of a more general EFT whose Lagrangian also includes operators of higher dimension obtained from the projection of a given UV theory onto the space of fields of the SM. The SM Effective Field Theory (SMEFT) realizes this idea by extending the Lagrangian of the SM by a complete set of Lorentz and (SM-)gauge invariant effective interactions of increasing mass dimensionality, built in terms of SM fields. These interactions do, in general, break the SM accidental symmetries but could also satisfy additional discrete and global symmetry properties, depending on the kind of NP they encode (see Falkowski (2023); Isidori et al. (2024) for a general introduction and Aebischer et al. (2025) for a state-of-the-art review of the SMEFT and its studies).
@@
Different UV extensions of the SM project differently on the effective interactions of the SMEFT Lagrangian below the UV scale $\Lambda$ de Blas et al. (2018); Guedes et al. (2023); Guedes and Olgoso (2024), such that constraining the effect of SMEFT interactions with available data could discriminate between classes of UV models and guide future BSM explorations. Most notably, different flavour symmetries assumed at the UV scale $\Lambda$ modify the number and structure of SMEFT interactions present below $\Lambda$ and directly affect a broad spectrum of observables. Several studies Aoude et al. (2020); Faroughy et al. (2020); Bißmann et al. (2021); Bruggisser et al. (2021, 2023); Bellafronte et al. (2023); Grunwald et al. (2023); Greljo and Palavrić (2023); Allwicher et al. (2023a); Bartocci et al. (2024); Mantani and Sanz (2025); Kala et al. (2025) have investigated the impact of different UV flavour assumptions and emphasised the need to move beyond the initial flavour-blind approach. A popular first step beyond assuming complete invariance under family rotation for the five kinds of fermions present in the SM (the so-called $U(3)^{5}$ flavour symmetry of the two left-handed doublets and three right-handed singlets of the SM gauge group) consists in singling out the third generation and still assuming invariance under rotation within the first two families (the so-called $U(2)^{5}$ flavour symmetry assumption). This choice is motivated by the scale hierarchy present among fermions and naturally appears in several BSM models. In our study we will consider both cases and discuss pros and cons of both assumptions in current fits. We emphasize that when specifying the flavour symmetry we only refer to the NP contribution encoded in the Wilson coefficients of operators of $d>4$. Typical examples of such NP models are given by heavy gauge bosons coupling universally to all flavours ($U(3)^{5}$, see e.g. ref. Salvioni et al. (2009)) or coupling differently to the third generation ($U(2)^{5}$). Since the $U(2)^{5}$ symmetry describes to a good approximation the quark flavour structure of the SM, one could also assume that the flavour invariance involves also the renormalizable part of the Lagrangian, with suitable breaking terms arising in both the $d\leq 4$ and $d>4$ terms in the SMEFT, as discussed in detail in ref. Faroughy et al. (2020) and typically happening in models seeking to explain the SM flavour hierarchy (see for example ref. Barbieri and Isidori (2024)).
@@
Overall, the great advantage of employing an effective field theory framework is that we can implement in a very general and systematic way a bottom-up approach to the exploration of some unknown UV theory using our knowledge of the physics we can currently test. Had patterns to emerge (e.g. enhanced or suppressed SMEFT coefficients, correlation and flat directions in SMEFT space, etc.) they could help in selecting classes of models whose investigation could then become more specific. The effective field theory framework is also useful in this regard. Using a top-down approach, one can easily re-interpret the SMEFT results in terms of specific models by using the matching between a given UV scenario and the effective theory. In the absence of unambiguous signs of BSM physics, the SMEFT general approach, although less informative, can indeed provide a very powerful exploratory tool. With this respect, we now have at our disposal for the first time a very broad spectrum of sufficiently accurate experimental measurements that can be simultaneously used to constrain the effective interactions of the SMEFT. This has indeed become a major effort in particle physics phenomenology and various comprehensive studies based on fits of the SMEFT that include a broad spectrum of observables have appeared in the recent literature Ellis et al. (2021); Ethier et al. (2021); Garosi et al. (2023); Allwicher et al. (2023a); Bartocci et al. (2024); Celada et al. (2024); Aad and others (2024c); Bartocci et al. (2025); Chekhovsky and others (2025); ter Hoeve et al. (2025).
@@
We will be working in the SMEFT between the UV scale $\Lambda$ and the EW scale ($\mu_{W}$), while at $\mu_{W}$ we will transition to the Low Energy Effective Theory (LEFT) Jenkins et al. (2018b) where all the heavy SM degrees of freedom ($W,Z,H,t$) are integrated out. To fully benefit of the EFT formalism and in order to account for impactful quantum effects, we will consider the scale dependence of the SMEFT and LEFT Wilson coefficients using leading-order RGE, with tree-level matching conditions at the EW scale $\mu_{W}$. All fits are performed in the open-source HEPfit framework J. de Blas et al. (2020); 1 using for each observable state-of-the-art theoretical predictions and the most recent experimental measurements.
The outcome of the fits presented in this paper will determine the ranges of SM parameters and SMEFT Wilson coefficients at the UV scale $\Lambda$ that are compatible with the selected broad set of experimental measurements. We will present both fits in which only one coefficient at a time is active and fits in which all SMEFT coefficients are simultaneously active at the scale $\Lambda$. Single-coefficient fits can be used to extract information on the set of observables that most constrain a given coefficient and can also give preliminary indications on the main RGE effects, while the global fit fully represents the combined effect of using a broad spectrum of observables and considering the operator mixing induced by the RGE. At the same time, a global fit of SM parameters and SMEFT coefficients can be very sensitive to the parametric and theoretical uncertainties and thus less effective, given current accuracies, to constrain the theory at hand. This dilution of the sensitivity is less evident in the $U(3)^{5}$ scenario, but has a strong impact in the $U(2)^{5}$ case, calling for the inclusion of more observables and/or for higher theoretical and experimental precision on the currently included ones.
@@
The paper consists of this introductory section followed by three main core sections and two appendices. A final section summarises the main conclusions of our study and presents an outlook on future developments. Concerning the core sections, in Section 2 we review elements of the SMEFT relevant to our study with particular emphasis on the implementation of the two aforementioned flavour scenarios, namely $U(3)^{5}$ and $U(2)^{2}$, and a detailed discussion of how the RGE of the SMEFT coefficients and SM parameters is included in our study. Technical details on how these two scenarios are implemented in the SMEFT basis of operators considered in our study are provided in Appendix A. Section 3 is dedicated to a discussion of the fitting framework and procedure, including how the various sets of observables have been implemented in HEPfit. A complete set of results is illustrated and discussed in Section 4 while the full outcome of both individual and global fits, for all three UV scales $\Lambda=1,\,3,$ and 10 TeV, with and without RGE, are collected in a series of tables in Appendix B.
@@ 2.1 The Standard Model Effective Field Theory: formalism and assumptions
where ${\cal O}_{i}^{(d)}$ are Lorentz and (SM)-gauge invariant operators of canonical mass dimension $d$ built only of SM fields, $i$ collectively labels each operator according to its field composition and quantum numbers, and $\Lambda$ is the cutoff of the effective theory indicatively identified with the UV scale of NP. If lepton number is conserved by the NP, the leading-order effects appear at dimension 6. In this work, we truncate the $\mathcal{L}_{\mathrm{SMEFT}}$ to include only dimension-6 operators and adopt the Warsaw basis introduced in ref. Grzadkowski et al. (2010). We also neglect baryon number violating and CP-odd operators. This results in the set of operators listed in Table 1.
@@
Table 1: Dimension-6 operators in the Warsaw basis, adapted from ref. Grzadkowski et al. (2010). We collectively denote by $X$ the field strength tensors of the SM gauge group ($X=G^{\mu\nu},W^{\mu\nu},B^{\mu\nu}$), by $\psi$ a generic fermion field, and by $\phi$ the scalar field of the SM. Among the fermion fields, $q$ and $l$ denote $SU(2)_{L}$ left-handed doublets while $u,d,e$ $SU(2)_{L}$ right-handed singlets, with family indices specified by the superscript in square brackets.
Furthermore, we do not consider the operator $\mathcal{O}_{\phi}$ since it does not enter at leading order in any of the observables considered in this study. In fact, the main effect of this operator is to modify the Higgs-boson self-interactions, and can only be weakly constrained with current Higgs pair-production data Tumasyan and others (2022a); Aad and others (2024a). We also notice that while this operator can be generated from other bosonic interactions via renormalization (to be discussed in Section 2.2), it does not enter in the 1-loop anomalous dimension matrix entries for any other SMEFT interaction, i.e. no other operator is generated from a non-zero value of $C_{\phi}$ at the UV scale.
@@
To define the $U(2)^{5}$ limit, it is necessary to specify a direction in flavour space that identifies the third generation. In this work, we choose the direction of the third family to be the one defined by the Yukawa couplings at the NP scale $\Lambda$, and consider the two possibilities where the third family is defined: 1) in the basis of diagonal up-type quark Yukawa couplings (denoted as UP basis in the following), or 2) in the basis of diagonal down-type quark Yukawa couplings (denoted as DOWN basis in the following). It is important to emphasise that these two classes of scenarios do not correspond to a mere change of basis, but rather reflect different assumptions about the flavour structure of the underlying UV physics. Notice that such a flavour misalignment arises exclusively in the left-handed quark sector — see the discussion in Silvestrini and Valli (2019) — while right-handed quark Yukawa couplings, as well as the full charged-lepton Yukawa couplings, can always be chosen flavour diagonal.
Having specified the basis, the $U(2)^{5}$ limit is described by a total of 124 (123 if not considering $C_{\phi}$) independent Wilson coefficients Faroughy et al. (2020). As in the $U(3)^{5}$ case, the relations between the $U(2)^{5}$ Wilson coefficients and those of the general SMEFT are presented in Appendix A.2, which also sets the notation for the presentation of the $U(2)^{5}$-symmetric results, following ref. Allwicher et al. (2023a).
@@ 2.2 Renormalization group evolution in the SMEFT
Assuming that the SMEFT coefficients are generated by BSM physics at the UV scale $\Lambda$, the Wilson coefficients at the EW scale $\mu_{W}$ are obtained by running them from the UV scale $\Lambda$ to the EW scale $\mu_{W}$ using the leading order (LO) RG equations derived in refs. Jenkins et al. (2013, 2014); Alonso et al. (2014). Lacking the next-to-leading SMEFT anomalous dimension, we can control only log-enhanced higher-order corrections to SMEFT contributions. Hence, we use tree-level matching conditions and matrix elements, and compute all SMEFT Wilson coefficients at a common scale $\mu_{W}$ that for convenience we choose to be $\mu_{W}=M_{W}$. For all high-$p_{T}$ observables (Higgs, top, and Drell-Yan), we use the same scale $\mu_{W}$, since the difference between the relevant physical scale of those processes and $M_{W}$ is such that no large logarithms arise, except for a few regions of phase space which currently still have large experimental uncertainties.
@@
One could be tempted to apply the same procedure to quark masses and CKM angles Descotes-Genon et al. (2019). However, one of the symmetry assumptions we are using, namely exact $U(2)^{5}$ for $d=6$ SMEFT operators, requires specifying the structure of the SM Yukawa couplings at the scale $\Lambda$, and the relations between quark masses, CKM angles, and Yukawa couplings are modified by SMEFT contributions. Thus, one cannot just trade the experimental values of quark masses and CKM angles for the corresponding SM parameters at the EW scale, and keeping the SMEFT contributions in the Yukawa matrices RGE is essential. [^1] To our knowledge, these corrections have been neglected in all the SMEFT fits that have considered the RGE evolution of the SMEFT coefficients so far. To impose the required flavour structure, we choose as parameters in the quark flavour sector the Yukawa couplings at the scale $\Lambda$, which are then evolved to the EW scale including the effect of the SMEFT in the running. Once evolved to the EW scale, the Yukawa couplings are used to compute the quark masses and CKM angles, including SMEFT contributions, which are then compared to the experimental values. Concerning the leptons, since we are assuming diagonal Yukawa couplings for the charged leptons and individual lepton numbers are preserved, the flavour orientation at the scale $\Lambda$ is irrelevant and, as stated above, we can trade the lepton Yukawa couplings at the EW scale for the corresponding masses.
The RG evolution from the scale $\Lambda$ to $\mu_{W}$ is performed within the HEPfit framework using the RGEsolver library Di Noi and Silvestrini (2023). The SMEFT Lagrangian at the scale $\mu_{W}$ is then expanded around the minimum of the scalar potential, and the quark mass matrices are diagonalised through the biunitary transformations
@@
The SMEFT operators involving quark fields are then transformed to the mass eigenstate basis using the matrices $V_{u,d}^{L,R}$ after removing unphysical phases. The unitary CKM matrix is defined as $V_{\mathrm{CKM}}\equiv V_{u}^{L\dagger}V_{d}^{L}$. [^2] Note that the determination of the CKM matrix elements from the experimental data is affected by additional corrections from SMEFT operators contributing to the relevant CKM observables. All these potential new physics effects are included in our calculations. The SMEFT coefficients in the mass eigenstate basis are then used to compute the observables at the scale $\mu_{W}$. [^3] Under our flavour assumptions, the flavour rotation is expected to have a negligible effect for the calculation of flavour-diagonal observables such as EW, Higgs boson, and top quark observables. Most observables in the flavour sector involve lower energies, and are therefore computed in the LEFT, where the massive gauge bosons, the Higgs field and the top quark are integrated out. The matching between the two EFT’s, the SMEFT and the LEFT, is implemented according to ref. Jenkins et al. (2018b). For each flavour observable, as detailed in Section 3.6, the LEFT coefficients are evolved to the relevant energy scale using the LEFT RG equations.